# Gaussian Process-Based Prediction and Control of Hammerstein-Wiener Systems

### v1

@@ Abstract
This work investigates data-driven prediction and control of Hammerstein-Wiener systems using physics-informed Gaussian process models. Data-driven prediction algorithms have been developed for structured nonlinear systems based on Willems’ fundamental lemma. However, existing frameworks cannot treat output nonlinearities and require a dictionary of basis functions for Hammerstein systems. In this work, an implicit predictor structure is considered, leveraging the multi-step-ahead ARX structure for the linear part of the model. This implicit function is learned by Gaussian process regression with kernel functions designed from Gaussian process priors for the nonlinearities. The linear model parameters are estimated as hyperparameters by assuming a stable spline hyperprior. The implicit Gaussian process model provides explicit output prediction by optimizing selected optimality criteria. The model is also applied to receding horizon control with the expected control cost and chance constraint satisfaction guarantee. Numerical results demonstrate that the proposed prediction and control algorithms are superior to black-box Gaussian process models.
Index Terms:
@@
Data-driven control, Gaussian processes, nonlinear systems.
I Introduction
@@
In recent years, the problem of data-driven prediction has drawn significant attention due to the emergence of more complex systems and the availability of big data. This problem aims to obtain nonparametric predictors directly from collected input-output data using minimal model knowledge of the system. In this way, model-based predictive controllers can be readily adapted into data-driven ones by employing data-driven predictors. This work investigates data-driven prediction and control of Hammerstein-Wiener systems. Hammerstein-Wiener systems are nonlinear block-oriented models with static input and output nonlinearities, which have a wide range of applications (see [1] and references therein).
@@
Data-driven prediction can be seen as finding a mapping from inputs and initial conditions to outputs, where any function approximators can be applied to learn the mapping. In this regard, a basis function decomposition approach is used in [2]. Gaussian process/kernel regression has been widely used in system identification to estimate the model [3, 4] by designing kernel structures for stable nonlinear systems inspired by system theory [5]. Kernel-based identification for Hammerstein and Wiener systems is investigated separately in [6, 7] for one-step-ahead prediction. This idea of using system-theoretic knowledge such as model structure in learning is sometimes known as physics-informed approaches [8]. Predictions from the Gaussian process model can then be used in receding horizon control. This approach is referred to as Gaussian process model predictive control [9, 10, 11].
@@
On the other hand, following the seminal result in [12], known as Willems’ fundamental lemma, data-driven prediction of linear systems has been widely studied by finding suitable linear combinations of collected data, subject to input and initial condition constraints [13]. The predictor can be reformulated as a multi-step-ahead ARX structure. Applying the predictor to receding horizon control leads to data-driven or data-enabled predictive control, for which multiple successful algorithms have been proposed, e.g., [14, 15, 16, 17, 18].
@@
Extensions of these linear algorithms to nonlinear systems have been analyzed for specific nonlinear model classes, including Hammerstein systems [19, 20], second-order Volterra systems [21], feedback linearizable systems [22], bilinear systems [23, 24], and systems with finite-dimensional Koopman embedding [25]. These extensions generalize Willems’ fundamental lemma by applying generalized signal embedding to the signal matrix and the persistency of excitation condition. This is possibly achieved by assuming a known dictionary of basis functions. However, for Hammerstein-Wiener systems, existing works [19, 24] address only the input nonlinearity (the Hammerstein part) and require prior knowledge of basis functions for the nonlinearity, which is often impractical.
In this work, data-driven prediction of Hammerstein-Wiener systems is first formulated as an implicit function learning problem by considering the multi-step-ahead ARX structure for the linear part. Instead of considering a basis function decomposition as in [19, 2, 20], Gaussian process priors are assumed for the nonlinearities and the kernel functions for the implicit prediction model are derived. The implicit Gaussian process model [26] then provides posterior predictions by optimizing selected optimality criteria, such as maximum likelihood. The linear model parameters act as hyperparameters in the Gaussian process model, which are estimated by enforcing a stable spline hyperprior to avoid overfitting [3, 4]. The nonlinearities can also be recovered separately using additional Gaussian process regression procedures. The implicit Gaussian process prediction is applied to receding horizon control by considering the expected control cost and guaranteeing output chance constraint satisfaction. Finally, the predictor and the control algorithm are verified in numerical examples and demonstrate better performance than the linear predictor and the black-box Gaussian process predictor with a nonlinear ARX model.
@@
The remainder of the paper is organized as follows. Section II introduces the problem and presents the implicit predictor structure. Section III proposes the data-driven prediction algorithm by Gaussian process regression. Section IV applies the data-driven predictor to receding horizon control, followed by numerical examples in Section V. Section VI concludes the paper.
@@
Notation. The expected value and covariance of a random variable are denoted by $\mathbb{E}[\cdot]$ and $\mathrm{cov}\left(\cdot\right)$, respectively. The symbol $\mathrm{Pr}\left(\cdot\right)$ indicates the probability of a random event. For a sequence of matrices $X_{1},\dots,X_{n}$, we denote the row-wise and diagonal-wise concatenation by $\mathrm{col}\left(X_{1},\dots,X_{n}\right)$ and $\mathrm{blkdiag}\left(X_{1},\dots,X_{n}\right)$, respectively. Given a signal $x:\left\{1,\dots,N\right\}\to\mathbb{R}^{n}$, the Hankel matrix of depth $L$ is defined as
@@
and its trajectory from $i$ to $j$ is indicated by $(x_{k})_{k=i}^{j}=\mathrm{col}\left(x_{i},\dots,x_{j}\right)$. For a vector $x$, $\left\lVert x\right\rVert_{P}$ denotes the weighted $l_{2}$-norm $(x^{\top}Px)^{\frac{1}{2}}$. The trace, log-determinant, and column vector containing the diagonal elements of a square matrix are indicated by $\mathrm{tr}\left(\cdot\right)$, $\mathrm{logdet}\left(\cdot\right)$, and $\mathrm{diag}\left(\cdot\right)$, respectively. The vectorization of a matrix is denoted by $\mathrm{vec}\left(\cdot\right)$. The symbols $\otimes$ and denote the Kronecker product and the Moore–Penrose inverse, respectively.
II Problem Statement and Background
$G(q)$$\psi(\cdot)$$\phi^{-1}(\cdot)$+$u_{k}$$\bar{u}_{k}$$\bar{y}_{k,0}$$e_{k}$$\bar{y}_{k}$$y_{k}$
@@ II Problem Statement and Background
Consider a Hammerstein-Wiener system shown in Fig. 1, given by
@@
where $x_{k}\in\mathbb{R}^{n_{x}}$, $u_{k}\in\mathbb{R}^{n_{u}}$, $y_{k}\in\mathbb{R}^{n_{y}}$ are the states, inputs, and outputs, respectively. The linear part of the system is denoted by $G(q)$, where $\bar{u}_{k}\in\mathbb{R}^{n_{u}}$, $\bar{y}_{k,0}\in\mathbb{R}^{n_{y}}$, $\bar{y}_{k}\in\mathbb{R}^{n_{y}}$, $e_{k}\in\mathbb{R}^{n_{y}}$ are the inputs, noise-free outputs, outputs, and output noise of $G(q)$, respectively. Functions $\psi(\cdot):\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{u}}$, $\phi^{-1}(\cdot):\mathbb{R}^{n_{y}}\rightarrow\mathbb{R}^{n_{y}}$ denote the input and output nonlinearities, respectively.
@@
1) The output noise $e_{k}\in\mathbb{R}^{n_{y}}$ is zero-mean i.i.d. Gaussian with covariance $\sigma^{2}\mathbb{I}$, 2) the linear part of the system $G(q)$ is stable, and 3) $\phi^{-1}(\cdot)$ is bijective.
Remark 1
The model is ambiguous up to scaling factors, i.e., the model triplet $\left(\psi(\cdot),G(q),\phi^{-1}(\cdot)\right)$ is equivalent to $\left(a\psi(\cdot),\frac{1}{ab}G(q),b\phi^{-1}(\cdot)\right)$, $a,b\in\mathbb{R}$. Thus, the nonlinearities $\psi(\cdot)$ and $\phi(\cdot)$ are normalized during learning in later sections.
@@
In this work, we are interested in characterizing probable length-$L$ input-output trajectory of (1) without model knowledge of $\psi(\cdot)$, $G(q)$, and $\phi^{-1}(\cdot)$, but from collected input-output trajectory $\left(\mathbf{u}^{d},\mathbf{y}^{d}\right):=\left(u_{k}^{d},y_{k}^{d}\right)_{k=1}^{N}$. This characterization is then used to design a receding horizon control algorithm.
II-A Data-Driven Prediction as Linear Systems
@@
This problem has been widely studied for linear systems, i.e., when $\psi(\cdot)$ and $\phi^{-1}(\cdot)$ are known. One successful approach recently relies on Willems’ fundamental lemma [12], skipping the explicit step of system identification. In detail, for $\mathbf{u}=(u_{k})_{k=1}^{L}$ and $\mathbf{y}=(y_{k})_{k=1}^{L}$, define $\Psi(\mathbf{u}):=\mathrm{col}\left(\psi(u_{1}),\dots,\psi(u_{L})\right)$ and $\Phi(\mathbf{y}):=\mathrm{col}\left(\phi(y_{1}),\dots,\phi(y_{L})\right)$. We have the following lemma.
Lemma 1
@@
Suppose $\left(\mathbf{u}^{d},\mathbf{y}^{d}\right):=\left(u_{k}^{d},y_{k}^{d}\right)_{k=1}^{N}$ is a noise-free trajectory of the Hammerstein-Wiener system (1). Let $\bar{\mathbf{u}}^{d}:=\Psi(\mathbf{u}^{d})$, $\bar{\mathbf{y}}^{d}:=\Phi(\mathbf{y}^{d})$. If $\bar{\mathbf{u}}^{d}$ is persistently exciting of order $(L+n_{x})$, i.e., $\mathcal{H}_{L+n_{x}}(\bar{\mathbf{u}}^{d})$ has full row rank, and $G(q)$ is controllable, $(\mathbf{u},\mathbf{y}):=\left(u_{k},y_{k}\right)_{k=1}^{L}$ is a trajectory of (1) iff there exists $g\in\mathbb{R}^{M}$, $M=N-L+1$, such that $\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}}\right)g=\mathrm{col}\left(\bar{\mathbf{u}},\bar{\mathbf{y}}_{0}\right)$, where $H_{\bar{u}}:=\mathcal{H}_{L}(\bar{\mathbf{u}}^{d})$, $H_{\bar{y}}:=\mathcal{H}_{L}(\bar{\mathbf{y}}^{d})$, $\bar{\mathbf{u}}:=\Psi(\mathbf{u})$, and $\bar{\mathbf{y}}_{0}:=\Phi(\mathbf{y})-\mathbf{e}$ denotes the noise-free output of $G(q)$, where $\mathbf{e}:=\left(e_{k}\right)_{k=1}^{L}$.
Proof:
This is a direct extension of Willems’ fundamental lemma, originally proposed in [12, Theorem 1]. ∎
@@
Let $L=:L_{0}+L^{\prime}$, where $L_{0}$ is no smaller than the lag of $G(q)$. There exists a mapping from $\mathbf{u}$ and $\mathbf{y}_{\text{p}}:=\left(y_{k}\right)_{k=1}^{L_{0}}$ to $\mathbf{y}_{\text{f}}:=\left(y_{k}\right)_{k=L_{0}+1}^{L}$. Define $\bar{\mathbf{y}}_{0}=:\mathrm{col}\left(\bar{\mathbf{y}}_{\text{p},0},\bar{\mathbf{y}}_{\text{f},0}\right)$ and $H_{\bar{y}}=:\mathrm{col}\left(H_{\bar{y}\text{p}},H_{\bar{y}\text{f}}\right)$, where $\bar{\mathbf{y}}_{\text{p},0}\in\mathbb{R}^{n_{y}L_{0}}$, $\bar{\mathbf{y}}_{\text{f},0}\in\mathbb{R}^{n_{y}L^{\prime}}$, $H_{\bar{y}\text{p}}\in\mathbb{R}^{n_{y}L_{0}\times M}$, and $H_{\bar{y}\text{f}}\in\mathbb{R}^{n_{y}L^{\prime}\times M}$. The mapping can be calculated by $g=\mathrm{col}\left(H_{\bar{u}},H_{\bar{y}\text{p}}\right)^{\dagger}\mathrm{col}\left(\bar{\mathbf{u}},\bar{\mathbf{y}}_{\text{p},0}\right)$, $\bar{\mathbf{y}}_{\text{f},0}=H_{\bar{y}\text{f}}g$. This leads to the following corollary.
Corollary 1
@@
An input-output sequence $(\mathbf{u},\mathbf{y})$ is a trajectory of (1) iff it satisfies
$\mathbf{0}=\left[\Gamma_{1}\ \Gamma_{2}\ -\mathbb{I}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\text{p}}),\Phi(\mathbf{y}_{\text{f}})\right)+\left[-\Gamma_{2}\ \mathbb{I}\right]\mathbf{e}.$ (2)

### v2

# Data-Driven Prediction and Control of Hammerstein-Wiener Systems with Implicit Gaussian Processes

@@ Abstract
This work investigates data-driven prediction and control of Hammerstein-Wiener systems using physics-informed Gaussian process (GP) models that encode the block-oriented model structure. Data-driven prediction algorithms have been developed for structured nonlinear systems based on Willems’ fundamental lemma. However, existing frameworks do not apply to output nonlinearities in Wiener systems and rely on a finite-dimensional dictionary of basis functions for Hammerstein systems. In this work, an implicit predictor structure is considered, leveraging the linearity for the dynamical part of the model. This implicit function is learned by GP regression, utilizing carefully designed structured kernel functions from linear model parameters and GP priors for the nonlinearities. Virtual derivative points are added to the regression by expectation propagation to encode monotonicity information of the nonlinearities. The linear model parameters are estimated as hyperparameters by assuming a stable spline hyperprior. The implicit GP model provides explicit output prediction by optimizing selected optimality criteria. The implicit model is also applied to receding horizon control with the expected control cost and chance constraint satisfaction guarantee. Numerical results demonstrate that the proposed prediction and control algorithms are superior to black-box GP models without model structure knowledge.
keywords:
@@
Data-driven control, Hammerstein-Wiener systems, Gaussian processes, kernel methods, nonlinear systems.
thanks: This paper was not presented at any IFAC meeting. This work was supported by the Lower Saxony Ministry for Science and Culture within the program zukunft.niedersachsen. Corresponding author Mingzhou Yin. Tel. +49 511 762 18903. Fax +49 511 762 4536.
1 Introduction
@@
In recent years, the problem of data-driven prediction has drawn significant attention due to the emergence of more complex systems and the availability of big data. This problem aims to obtain nonparametric predictors directly from collected input-output data using minimal model knowledge of the system. In this way, model-based predictive controllers can be readily adapted into data-driven ones by employing data-driven predictors. This work investigates data-driven prediction and control of Hammerstein-Wiener (H-W) systems. H-W systems are nonlinear block-oriented models with static input and output nonlinearities [14], which are widely used to model, e.g., chemical processes [1], electromechanical systems [47], and ionospheric dynamics [30]. More generally, the input and output nonlinearities can be used to model actuator and sensor nonlinearities, respectively.
@@
Data-driven prediction can be seen as finding a mapping from inputs and initial conditions to outputs, where any function approximator can be applied to learn the mapping. In this regard, a basis function decomposition approach is used in [22]. Gaussian process (GP) or kernel regression has been widely used in system identification to estimate the model [31, 32] by designing kernel structures for stable nonlinear systems inspired by system theory [33]. Kernel-based identification for Hammerstein and Wiener systems is investigated separately in [37, 38] for one-step-ahead prediction. This idea of using system-theoretic knowledge (such as model structure) in learning is sometimes known as physics-informed approaches [11]. Predictions from the GP model can then be used in receding horizon control. This approach is referred to as Gaussian process model predictive control (GP-MPC) [19, 15, 7, 24, 40, 34]. However, existing GP-MPC schemes employ either a fully black-box model without any model structure or an additive model structure, consisting of a known nominal model plus black-box dynamics, neither of which can encode the H-W structure to obtain a physics-informed model. In addition, these schemes only learn a one-step-ahead predictor and require uncertainty propagation over the prediction horizon, which is challenging and often requires approximation [15, 40].
@@
On the other hand, following the seminal result in [43], known as Willems’ fundamental lemma (WFL), data-driven prediction of linear systems has been widely studied by finding suitable linear combinations of persistently exciting data, subject to input and initial condition constraints [26]. The predictor can be reformulated as a multi-step-ahead autoregressive with extra input (ARX) structure. Applying the predictor to receding horizon control leads to data-driven (or data-enabled) predictive control (DDPC), for which multiple successful algorithms have been proposed, e.g., [10, 6, 45, 8, 2, 9].
@@
Extensions of these linear algorithms to nonlinear systems have been analyzed for specific nonlinear model classes, including Hammerstein systems [5, 25, 29], second-order Volterra systems [39], feedback linearizable systems [3, 29], bilinear systems [46, 25], and systems with finite-dimensional Koopman embedding [42]. These extensions generalize WFL by applying generalized signal embedding to the signal matrix. Some also apply a kernelized version of WFL by considering the inner product of the signal embedding using the “kernel trick” [29, 23, 16]. However, for H-W systems, existing works [5, 25, 29] address only the input nonlinearity (the Hammerstein part) but not the output nonlinearity (the Wiener part) [^1] Wiener systems are discussed in [5], but an extension of WFL that enables data-driven prediction is not obtained for this system class, cf. Section 2.2 below.. In addition, the signal embedding or its kernel reformulation requires a known finite-dimensional dictionary of basis functions to satisfy the persistency of excitation condition. Such information can be impractical to obtain. The lifted signal space also makes it difficult to satisfy the persistency of excitation condition, requiring a long input sequence with tailored design [4].
In this work, data-driven prediction of H-W systems is first formulated as an implicit function learning problem by considering the multi-step-ahead ARX structure for the linear part. GP priors are assumed for the nonlinearities, which correspond to an infinite-dimensional reproducing kernel Hilbert space (RKHS), where universal approximation capability can be attained for common designs (such as squared exponential kernels) without any knowledge about the nonlinearities [28]. Then, the structured kernel function for the implicit prediction model is derived together with the linear model parameters. Compared with an explicit black-box GP model with general kernel functions, the structured kernel design provides a physics-informed design by considering a smaller function space for the predictor that is compatible with the H-W model structure.
To retain the well-definedness of the output nonlinearities in the implicit GP model, virtual derivative points are added to promote monotonic output nonlinearities by expectation propagation [35, 36]. The implicit GP model [27] provides explicit posterior predictions by optimizing selected optimality criteria, such as maximum likelihood. The linear model parameters act as hyperparameters in the proposed GP model, which are estimated by enforcing a stable spline hyperprior and solving a joint maximum-a-posteriori/maximum-likelihood (JMAP-ML) problem to avoid overfitting [31, 32]. The nonlinearities can also be recovered separately using additional GP regression procedures.
The implicit GP predictor is applied to receding horizon control by considering the expected control cost and guaranteeing output chance constraint satisfaction. The proposed predictor provides multi-step-ahead predictions by nature, circumventing the difficult problem of uncertainty propagation in standard GP-MPC with one-step-ahead prediction. Finally, the predictor and the control algorithm are verified in numerical examples and demonstrate better performance than the black-box GP predictor with a general nonlinear ARX model.
@@
The remainder of the paper is organized as follows. Section 2 introduces the problem and presents the implicit predictor structure. Section 3 proposes the data-driven prediction algorithm by implicit GP regression. Section 4 applies the data-driven predictor to receding horizon control, followed by numerical examples in Section 5. Section 6 concludes the paper.
@@
Notation. The expected value and covariance of a random variable are denoted by $\mathbb{E}[\cdot]$ and $\mathrm{cov}\left(\cdot\right)$, respectively. The symbol $\mathrm{Pr}\left(\cdot\right)$ indicates the probability of a random event. A GP is indicated by $\mathcal{GP}\left(m(\cdot),k(\cdot,\cdot)\right)$, where $m(\cdot)$ and $k(\cdot,\cdot)$ are the mean and kernel/covariance functions, respectively. For a sequence of matrices $X_{1},\dots,X_{n}$, we denote the row-wise and diagonal-wise concatenation by $\mathrm{col}\left(X_{1},\dots,X_{n}\right)$ and $\mathrm{blkdiag}\left(X_{1},\dots,X_{n}\right)$, respectively. For a vector $x$, $\left\lVert x\right\rVert_{P}$ denotes the weighted $l_{2}$-norm $(x^{\top}Px)^{\frac{1}{2}}$. The trace, determinant, log-determinant, and column vector containing the diagonal elements of a square matrix are indicated by $\mathrm{tr}\left(\cdot\right)$, $\mathrm{det}\left(\cdot\right)$, $\mathrm{logdet}\left(\cdot\right)$, and $\mathrm{diag}\left(\cdot\right)$, respectively. The vectorization, null space, and the $(i,j)$-th element of a matrix are denoted by $\mathrm{vec}\left(\cdot\right)$, $\mathrm{null}\left(\cdot\right)$, and $(\cdot)_{i,j}$, respectively. Given a signal $x:\left\{1,\dots,N\right\}\to\mathbb{R}^{n}$, the Hankel matrix of depth $L$ is defined as
@@
and its trajectory from $i$ to $j$ is indicated by $(x_{k})_{k=i}^{j}=\mathrm{col}\left(x_{i},\dots,x_{j}\right)$. The symbols $\otimes$ and denote the Kronecker product and the Moore–Penrose inverse, respectively.
[…]
@@
2 Problem Statement and Background
Consider a H-W system shown in Fig. 1, given by
@@ 2 Problem Statement and Background
where $x_{k}\in\mathbb{R}^{n_{x}}$, $u_{k}\in\mathbb{R}^{n_{u}}$, $y_{k}\in\mathbb{R}^{n_{y}}$ are the states, inputs, and outputs, respectively. The linear part of the system is denoted by $G(q)$, where $\bar{u}_{k}\in\mathbb{R}^{n_{u}}$, $\bar{y}_{k,0}\in\mathbb{R}^{n_{y}}$, $\bar{y}_{k}\in\mathbb{R}^{n_{y}}$, $e_{k}\in\mathbb{R}^{n_{y}}$ are the inputs, noise-free outputs, outputs, and output noise of $G(q)$, respectively. We also define $y_{k,0}=\phi^{-1}(\bar{y}_{k,0})$ as the noise-free nonlinear output. The functions $\psi(\cdot):\mathcal{U}\rightarrow\bar{\mathcal{U}}$, $\phi^{-1}(\cdot):\bar{\mathcal{Y}}\rightarrow\mathcal{Y}$, where $\mathcal{U},\bar{\mathcal{U}}\subseteq\mathbb{R}^{n_{u}}$ and $\mathcal{Y},\bar{\mathcal{Y}}\subseteq\mathbb{R}^{n_{y}}$ denote the input and output nonlinearities, respectively.
@@
1) The output noise $e_{k}\in\mathbb{R}^{n_{y}}$ is zero-mean i.i.d. Gaussian with covariance $\sigma^{2}\mathbb{I}$, 2) the linear part of the system $G(q)$ is stable, and 3) $\phi(\cdot)$ is monotonically increasing, i.e., the Jacobian of $\phi(y)$ is positive-definite for all $y$.
Assumption 1.3 guarantees that $\phi^{-1}(\cdot)$ is well-defined and is reasonable when $\phi^{-1}(\cdot)$ depicts, for example, sensor nonlinearity.
@@
In this work, we are interested in characterizing probable length-$L$ input-output trajectories of (1) from a collected input-output trajectory $\left(\mathbf{u}^{d},\mathbf{y}^{d}\right):=\left(u_{k}^{d},y_{k}^{d}\right)_{k=1}^{N}$ by incorporating the H-W model structure but not model knowledge of the components $\psi(\cdot)$, $G(q)$, and $\phi^{-1}(\cdot)$. This characterization is then used to design a receding horizon control algorithm.
2.1 Data-Driven Prediction for Linear Systems
@@
This problem has been widely studied for linear systems, i.e., when $\psi(\cdot)$ and $\phi^{-1}(\cdot)$ are known. One successful approach recently relies on WFL [43], skipping the explicit step of system identification. In detail, for $\mathbf{u}=(u_{k})_{k=1}^{L}$ and $\mathbf{y}=(y_{k})_{k=1}^{L}$, define $\Psi(\mathbf{u}):=\mathrm{col}\left(\psi(u_{1}),\dots,\psi(u_{L})\right)$ and $\Phi(\mathbf{y}):=\mathrm{col}\left(\phi(y_{1}),\dots,\phi(y_{L})\right)$. If noise-free trajectory data are available, we have the following lemma, which is a direct application of WFL, originally proposed in [43, Theorem 1], on the linear part of the H-W system.
Lemma 1.
@@
Suppose $\left(\mathbf{u}^{d},\mathbf{y}^{d}\right):=\left(u_{k}^{d},y_{k}^{d}\right)_{k=1}^{N}$ is a noise-free trajectory of the H-W system (1). Let $\bar{\mathbf{u}}^{d}:=\Psi(\mathbf{u}^{d})$, $\bar{\mathbf{y}}^{d}:=\Phi(\mathbf{y}^{d})$. If $\bar{\mathbf{u}}^{d}$ is persistently exciting of order $(L+n_{x})$, i.e., $\mathcal{H}_{L+n_{x}}(\bar{\mathbf{u}}^{d})$ has full row rank, and $G(q)$ is controllable, an input-output sequence $(\mathbf{u},\mathbf{y}):=\left(u_{k},y_{k}\right)_{k=1}^{L}$ is a trajectory of (1) iff there exists $g\in\mathbb{R}^{M}$, M=NL+1M=N-L+1, such that
$\begin{bmatrix}H_{\bar{u}}\\ H_{\bar{y}}\end{bmatrix}g=\begin{bmatrix}\Psi(\mathbf{u})\\ \Phi(\mathbf{y})-\mathbf{e}\end{bmatrix}$,
where $H_{\bar{u}}:=\mathcal{H}_{L}(\bar{\mathbf{u}}^{d})$, $H_{\bar{y}}:=\mathcal{H}_{L}(\bar{\mathbf{y}}^{d})$, and $\mathbf{e}:=\left(e_{k}\right)_{k=1}^{L}$.
Lemma 1 is not directly applicable to our problem since we have neither noise-free trajectory data, nor the knowledge of the nonlinearities. Nevertheless, it is instructive for providing a suitable structure to learn the predictor, as detailed below.
Let $L=:L_{0}+L^{\prime}$, where $L_{0}$ is no smaller than the lag of $G(q)$. Define the following partitions:
$\mathbf{y}=:\begin{bmatrix}\mathbf{y}_{\mathrm{p}}\\ \mathbf{y}_{\mathrm{f}}\end{bmatrix},\ \mathbf{e}=:\begin{bmatrix}\mathbf{e}_{\mathrm{p}}\\ \mathbf{e}_{\mathrm{f}}\end{bmatrix},\ H_{\bar{y}}=:\begin{bmatrix}H_{\bar{y}\mathrm{p}}\\ H_{\bar{y}\mathrm{f}}\end{bmatrix}$,
@@
where $\mathbf{y}_{\mathrm{p}},\mathbf{e}_{\mathrm{p}}\in\mathbb{R}^{n_{y}L_{0}}$, $\mathbf{y}_{\mathrm{f}},\mathbf{e}_{\mathrm{f}}\in\mathbb{R}^{n_{y}L^{\prime}}$, $H_{\bar{y}\mathrm{p}}\in\mathbb{R}^{n_{y}L_{0}\times M}$, and $H_{\bar{y}\mathrm{f}}\in\mathbb{R}^{n_{y}L^{\prime}\times M}$. Then, there exists a mapping from $\Psi(\mathbf{u})$ and $\Phi(\mathbf{y}_{\mathrm{p}})$ to $\Phi(\mathbf{y}_{\mathrm{f}})$, which is characterized by the following corollary.
Corollary 1.
@@
Under the same conditions as Lemma 1, an input-output sequence $(\mathbf{u},\mathbf{y})$ is a trajectory of (1) iff it satisfies
$\mathbf{0}=\left[\Gamma_{1}\ \Gamma_{2}\ -\mathbb{I}\right]\mathrm{col}\left(\Psi(\mathbf{u}),\Phi(\mathbf{y}_{\mathrm{p}}),\Phi(\mathbf{y}_{\mathrm{f}})\right)-\left[\Gamma_{2}\ -\mathbb{I}\right]\mathbf{e}.$ (2)